Evaluate the project having following cash flow with both types of B/C ratio using by PW method
8
Initial investment =Rs. 1,000,000
Salvage value =Rs. 150,000
Annual revenue =Rs. 190,000
Annual expenses =Rs. 50,000
MARR = 12% per year
Useful life of project = 12 Years
1
Expert's answer
2020-05-27T09:21:41-0400

"PW_B=190,000(P/A,12\%,12)+150,000(P/F,12\%,12)"

"=190,000×\frac{(1+0.12)^12-1}{0.12×(1+0.12)^12}+150,000×(1+0.12)^12=1,215,432"

"PW_C=1,000,000+50,000(P/A,12\%,12)"

"=1,000,000+50,000×\frac{(1+0.12)^12-1}{0.12×(1+0.12)^12}=1,309,719"

"B/C=\frac{PW_B}{PW_C}=\frac{1,215,432}{1,309,713}=0.92"